Circular No. 7704

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001du IN NGC 1365 S. J. Smartt, Institute of Astronomy, University of Cambridge; D.

Kilkenny, South African Astronomical Observatory (SAAO); and P. Meikle, Imperial College, London, report that a spectrum taken at the SAAO 1.9-m telescope (+ Cassegrain Spectrograph) on Sept. 2.05 UT indicates that SN 2001du (IAUC 7690, 7692) is a type-II supernova: “The spectrum (340– 530 nm, resolution 0.3 nm) has a very blue continuum, indicating an early epoch and eruption likely to have been just before Aug. 24 (cf. IAUC 7690). There is a strong Hβ P-Cyg profile, and the Hγ, Hδ, and Hϵ lines are seen in absorption. The expansion velocity determined from the trough of the Hβ absorption is ∼ 6000 km/s (assuming the NED recession velocity for NGC 1365 of 1636 km/s). This galaxy was well studied by the Hubble Space Telescope during the Cepheid-distance-scale Key Project, and has a distance of 18.3 ± 3.3 Mpc (distance modulus 31.31 ± 0.38); the site of SN 2001du falls on three pre-explosion images taken with WFPC2, which could possibly allow a progenitor to be identified. At the very least, limits will be set on the progenitor mass (as in Smartt et al. 2001, Ap.J. 556, L29). Hence, observations of this event at optical, radio, infrared, and x-ray wavelengths are strongly encouraged.”

L. Wang, Lawrence Berkeley National Laboratory; D. Baade, European Southern Observatory (ESO); C. Fransson, Stockholm Observatory; P. Hoeflich, University of Texas; P. Lundqvist, Stockholm Observatory; and J. C. Wheeler, University of Texas, write: “SN 2001du was observed with the ESO Very Large Telescope UT3 (+ FORS1) on Aug. 30 UT. The spectra (range 416–860 nm; spectral resolution 1.27 nm and 0.265 nm/pixel) confirm that this is a type-IIP supernova before maximum. Strong Hα, Hβ, and He I 587.6-nm lines were detected. The P-Cyg absorption trough of the Hα line is blueshifted by 10 700 km/s after correction for the recession velocity of the host galaxy. We have obtained high-quality spectropolarimetry data of SN 2001du (noise level per pixel ∼ 0.15 percent). The supernova is found to be polarized at a level of 0.29 percent with a flat spectrum characteristic of those caused by electron scattering.”

SUPERNOVA 2001ed IN NGC 706 Pre-discovery magnitudes of SN 2001ed (cf. IAUC 7703): Aug. 21 UT,

V = 14.7 ± 0.5 (M. Salvo and B. Schmidt, Mount Stromlo Observatory, from Southern All-Sky Supernova Survey images); 22.13, ∼ 14.5 (P. Antonini, Avignon, France, unfiltered CCD); 29.13, ∼ 15.0 (Antonini).

2001 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green